# Escaping the BLEU Trap: A Signal-Grounded Framework with Decoupled Semantic Guidance for EEG-to-Text Decoding

### v1

@@ 1 Introduction
To this end, we propose SemKey, a novel two-stage framework enforcing semantic guidance and strict signal grounding. In the first stage, the Parallel Multi-task Attribute Extraction stage is inspired by neuroscientific findings that high-level semantics form robust representations (Lerner et al., 2011). Therefore, we decouple the decoding process by explicitly predicting four attributes (sentiment, topic, length, and surprisal) to mitigate semantic bias and reduce repetitive templates. These serve as stable semantic anchors to steer the LLM, effectively breaking the cycle of bias and repetition. Then, in the second stage, the Multi-Perspective Active Retrieval Decoding stage, to mitigate signal neglect, we redesign the encoder-LLM interaction via a Q-K-V Injection mechanism. Recognizing that standard cross-attention often fails to override the LLM’s strong language modeling objective, we treat EEG embeddings as Keys/Values explicitly queried by multi-perspective semantic prompts obtained from the first stage. This mechanism transforms decoding from passive conditioning to active information retrieval, ensuring generation remains strictly grounded in neural input. Furthermore, based on our observation that previous baselines achieve high BLEU scores primarily through template memorization rather than genuine decoding, we advocate moving beyond this “BLEU trap”. We establish a comprehensive evaluation protocol using rigorous Retrieval and Distribution metrics. Experiments demonstrate that SemKey achieves state-of-the-art performance, outperforming the GLIM baseline by 42.7% in Self-BLEU, 54.4% in FD, and 36.4% in 24-Way Retrieval Accuracy.
@@ 3.1 Stage 1: Parallel Multi-Task Attribute Extraction
Based on (Liu et al., 2025), we employ a multi-task learning strategy to explicitly decouple high-level semantic labels. The EEG encoder transforms signals into latent embeddings, inheriting alignment and reconstruction tasks via a frozen Flan-T5 model (Chung et al., 2024). Concurrently, we introduce four auxiliary decoding heads to predict semantic attributes: Sentiment, Topic, Length, and Surprisal. These predictions serve as a robust semantic skeleton to guide the subsequent generation. As shown in Figure 2, this stage is structured into three key components: a prior-guided encoding stream, a parallel multi-task learning framework, and a unified training objective.
@@ 4.1 Main Performance Comparison
EEG-to-Text* 74.8% 43.6% 19.0% 8.1% 2.1% 0.1% 0.1% 0.00 100.0% 1.65
GLIM 83.3% 57.5% 29.7% 14.3% 1.2% 3.0% 8.1% 2.18 90.8% 0.57
SemKey(Ours) 86.1% 62.8% 36.7% 19.5% 2.6% 5.3% 22.6% 6.21 52.0% 0.26
Rel. Improv. +3.4% +9.2% +23.6% +36.4% +116.7% +76.7% +179.0% +184.9% +42.7% +54.4%
@@
First, regarding Alignment & Semantic Substance, SemKey consistently outperforms baselines across all retrieval settings ($N=2,4,10,24$). Notably, the performance advantage widens as task difficulty increases: while achieving a 3.4% relative gain in the 2-Way setting, this margin expands significantly to +23.6% in 10-Way and +36.4% in 24-Way scenarios. This trend indicates that our model is substantially more robust in distinguishing fine-grained semantic nuances against a large pool of distractors.
@@ 4.4 Ablation Studies
w/o Topic 58.3% 15.8% 12.7% 0.40
w/o Sentiment 58.1% 15.7% 15.5% 0.42
w/o Length 57.5% 15.3% 9.6% 0.67
w/o Surprisal 55.4% 14.0% 4.8% 0.63
SemKey (Full) 62.8% 19.5% 22.6% 0.26
@@ 1. LLM Tuning Strategy
LLM Full Fine-tune 61.1% 18.3% 22.2% 0.29
LLM + LoRA 61.7% 18.4% 21.3% 0.30
LLM Frozen (Ours) 62.8% 19.5% 22.6% 0.26
@@ 3. Stage 2 EEG Encoder
Encoder Fixed 58.2% 15.3% 12.2% 0.51
Encoder Trainable (Ours) 62.8% 19.5% 22.6% 0.26
@@ B.3 Dataset & Preprocessing
Table d: Ablation study on Spectral Whitening. Whitening significantly boosts performance, yielding up to 21.8% relative improvement in diversity (Dist-2) and 8.6% in retrieval accuracy (24-Way).
@@
   w/o Whitening    60.63%    17.94%    18.56    2.35%
   w/ Whitening (Ours)    62.77%    19.49%    22.60    2.61%
   Relative Improv.    +3.53%    +8.64%    +21.77%    +11.06%

### v2

@@ 1 Introduction
To this end, we propose SemKey, a novel two-stage framework enforcing semantic guidance and strict signal grounding. In the first stage, the Parallel Multi-task Attribute Extraction stage is inspired by neuroscientific findings that high-level semantics form robust representations (Lerner et al., 2011). Therefore, we decouple the decoding process by explicitly predicting four attributes (sentiment, topic, length, and surprisal) to mitigate semantic bias and reduce repetitive templates. These serve as stable semantic anchors to steer the LLM, effectively breaking the cycle of bias and repetition. Then, in the second stage, the Multi-Perspective Active Retrieval Decoding stage, to mitigate signal neglect, we redesign the encoder-LLM interaction via a Q-K-V Injection mechanism. Recognizing that standard cross-attention often fails to override the LLM’s strong language modeling objective, we treat EEG embeddings as Keys/Values explicitly queried by multi-perspective semantic prompts obtained from the first stage. This mechanism transforms decoding from passive conditioning to active information retrieval, ensuring generation remains strictly grounded in neural input. Furthermore, based on our observation that previous baselines achieve high BLEU scores primarily through template memorization rather than genuine decoding, we advocate moving beyond this “BLEU trap”. We establish a comprehensive evaluation protocol using rigorous Retrieval and Distribution metrics. Experiments demonstrate that SemKey achieves state-of-the-art performance, outperforming the GLIM baseline by 42.7% in Self-BLEU, 54.4% in FD, and 51.2% in 24-Way Retrieval Accuracy.
@@ 3.1 Stage 1: Parallel Multi-Task Attribute Extraction
Based on (Liu et al., 2025), we employ a multi-task learning strategy to explicitly decouple x-level semantic labels. The EEG encoder transforms signals into latent embeddings, inheriting alignment and reconstruction tasks via a frozen Flan-T5 model (Chung et al., 2024). Concurrently, we introduce four auxiliary decoding heads to predict semantic attributes: Sentiment, Topic, Length, and Surprisal. These predictions serve as a robust semantic skeleton to guide the subsequent generation. As shown in Figure 2, this stage is structured into three key components: a prior-guided encoding stream, a parallel multi-task learning framework, and a unified training objective.
@@ 4.1 Main Performance Comparison
EEG-to-Text* 50.0% 25.0% 10.0% 4.17% 2.1% 0.1% 0.1% 0.00 100.0% 1.65
GLIM 66.8% 42.4% 21.9% 10.7% 1.2% 3.0% 8.1% 2.18 90.8% 0.57
SemKey(Ours) 72.2% 50.4% 29.5% 16.2% 2.6% 5.3% 22.6% 6.21 52.0% 0.26
Rel. Improv. +8.0% +18.9% +34.4% +51.2% +116.7% +76.7% +179.0% +184.9% +42.7% +54.4%
@@
First, regarding Alignment & Semantic Substance, SemKey consistently outperforms baselines across all retrieval settings ($N=2,4,10,24$). Notably, the performance advantage widens as task difficulty increases: while achieving a 8.0% relative gain in the 2-Way setting, this margin expands significantly to +34.4% in 10-Way and +51.2% in 24-Way scenarios. This trend indicates that our model is substantially more robust in distinguishing fine-grained semantic nuances against a large pool of distractors.
@@ 4.4 Ablation Studies
w/o Topic 43.7% 12.3% 12.7% 0.40
w/o Sentiment 44.5% 12.1% 15.5% 0.42
w/o Length 43.5% 11.7% 9.6% 0.67
w/o Surprisal 40.7% 10.4% 4.8% 0.63
SemKey (Full) 50.4% 16.2% 22.6% 0.26
@@ 1. LLM Tuning Strategy
LLM Full Fine-tune 47.9% 14.8% 22.2% 0.29
LLM + LoRA 49.1% 14.8% 21.3% 0.30
LLM Frozen (Ours) 50.4% 16.2% 22.6% 0.26
@@ 3. Stage 2 EEG Encoder
Encoder Fixed 44.2% 11.6% 12.2% 0.51
Encoder Trainable (Ours) 50.4% 16.2% 22.6% 0.26
@@ B.3 Dataset & Preprocessing
Table d: Ablation study on Spectral Whitening. Whitening significantly boosts performance, yielding up to 21.8% relative improvement in diversity (Dist-2) and 11.41% in retrieval accuracy (24-Way).
@@
   w/o Whitening    48.00%    14.55%    18.56    2.35%
   w/ Whitening (Ours)    50.35%    16.21%    22.60    2.61%
   Relative Improv.    +4.90%    +11.41%    +21.77%    +11.06%